My daughter trying this math problem 5(3-7a)=-230 here's her work
-15-35a=-230
-35a=-215/-35
No answer cause she said it would b a decimal pls help so I can show her how to so her other problems tks

Answer by ewatrrr(24785)

 
Hi,
The Idea is to Isolate 'a' on one side of the Equation
5(3-7a)=-230 |Distributing the 5
15-35a = -230 |Add 35a to both sides
15 = -230 + 35a |add 230 to both sides
245 = 35a |Divide both sides by 35
7 = a